Question 3: Details of proposed project of Samex LLC are given below: Initial Investment: $120,000 Expected life: 4 years (zero scrap value) Estimated Annual Cash Flows: Year 1 $50,000 Year 2 $45,000 Year 3 $30,000 Year 4 $30,000 Estimated cost of capital is 10%

Calculate: a. Payback Period b. Accounting Rate of Return
Answer #1

Initial Investment CF0=120000,n=4 years.

Cash Inflows. CF1=50000,CF2=45000, CF3=30000,CF4=30000.

When uneven casf flow is given .Paaay back is calculated using cumulative cash inflows.

Year Cash In Cumulative cash in
1 50000 50000
2 45000 95000
3 30000 125000
4 30000 155000

Hence 95,000 is collecetd at the end of 2nd year.we require initial investment of 120000.

i.e 25,000(120000-95000) more is required.

In the 3rd year Cash inflow of 30000 is received .Hence payback fraction is 25,000/30000=0.833.

Hence payback period is 2.833 years.

Accounting Rate of return=Average income/Average Investment*100

Average income is 155000/4=38,750.

average investment=Net working capital+Salvage value+0.5*(Initial cost of machine-Salvage Value)

salvage value or scrap value=0 given

Hence ARR=(38,750/120000)*100

32.3%.

These are called as Traditional Capital Budgeting methods .These don't use discounting cash flow techniques hence Cost of capital is not used.